OpenStudy (renren):

Strep throat is caused by Streptococcus bacteria. The amount of bacteria multiplies at a rate of 10^13 per 10^3 hrs. What is the average growth of bacteria per hour??

OpenStudy (anonymous):

What are the options?

563blackghost (563blackghost):

Is this multiple choice?

563blackghost (563blackghost):

What I would do is divide them both....to find the unit rate...

OpenStudy (anonymous):

10^13/10^3 = your answer When dividing exponents w/ the same base, just subtract the powers ex: 10^5/10^2 = 10^3

563blackghost (563blackghost):

was gonna put down how to divide the exponents by subtracting but u beat me to it ^^ @robjoe

OpenStudy (anonymous):

Next time, @563blackghost :D (can u help w/ my stats question?)

OpenStudy (renren):

Thank you very much. My question was answered. Robjoe

OpenStudy (anonymous):

@renren no prob man!

OpenStudy (renren):

Yea mon

OpenStudy (renren):

The mass of planet mercury is 3.3*10^23 kilograms. The mass of planet Jupiter is 5.8*10^3 times the mass of planet Mercury. Planet Jupiter is 3.2*10^2 times the mass of planet Earth. Part A Using scientific notation, what is the mass of planet Jupiter? Answer in Kilogram Part B Using scientific notation, what is the mass of planet Earth? Answer in Kilogram